OpenStudy (anonymous):

Sending text messages from Katrina's cell phone costs $0.14 per text message. From last month, she has a credit on her bill for $7.00. How much more money will she need to pay for text messages this month if she has sent 184 messages? $20.26 $18.76 $16.76 $25.76

OpenStudy (anonymous):

price of 184messages=184*0.14=25.76 money that she have to top up=total price-credit in her bill=25.76-7=18.76 So answer is B

OpenStudy (anonymous):

thnk u